OpenStudy (anonymous):

Find a third–degree polynomial function such that f(0) = –24 and whose zeros are –1, 2, and –3. Using complete sentences, explain how you found it.

OpenStudy (anonymous):

\[f(x)=a(x+1)(x-2)(x+3)\] solve for a via \[f(0)=24=a\times 1\times (-2)\times 3\]